OpenStudy (leann01xxx):

Al goes 8 mph; Ben goes 22 mph in the opposite direction. How many hours has Ben traveled when they are 180 miles apart? A. 15 B. 6 C. 12.86

OpenStudy (mathstudent55):

\(speed = \dfrac{distance}{time} \) \(distance = speed \cdot time\) They both traveled the same time, t. Using the second equation, what distance did each one travel in t time?

OpenStudy (leann01xxx):

I have no idea. I failed this lesson of algebra.

OpenStudy (pawanyadav):

Ben velocity relative to All =22-(-8)mph=30mph Due opposite direction Now time taken =180÷30hour=6hour

OpenStudy (pawanyadav):

This is solved by Physics

OpenStudy (leann01xxx):

So then is it C?

OpenStudy (rhystic):

They both travelled for the same amount of time and reached a total distance of 180 miles. thus utilizing distance = speed*time you have distance = speed1*time+speed2*time let time =x and you have distance = speed1x + speed2x. or 180= 8x + 22x or 180 = 30x . solve for x and you have your time in hours.

OpenStudy (mathstudent55):

Al goes 8 mph d1 = 8x Ben goes 22 mph in the opposite direction d2 = 22x The total distance traveled is d1 + d2, but we are told that is 180 miles, so 8x + 22x = 180 30x = 180 x = 6 Answer: 6 hours